\begin{thebibliography}{99}
%
\bibitem{AAD2001}
A.A.~Abramov, A.~Aslanyan, and E.B.~Davies, 
\emph{Bounds on complex eigenvalues and resonances}, 
Journal of Physics A: Mathematical and General, \textbf{34} (2001), no. 1, 57.
%
\bibitem{B_F_R_V_V}
J.A.~Barcel\'o, L.~Fanelli, A.~Ruiz, M.C.~Vilela, and N.~Visciglia,
\emph{Resolvent and Strichartz estimates for elastic wave equations},
Appl. Math. Lett., \textbf{49} (2015), 33--41.
%
\bibitem{B_F-G_P-E_R_V}
J.A.~Barcel\'o, M.~Folch-Gabayet, S.~Pérez-Esteva, A.~Ruiz, and M.C.~Vilela,
\emph{Limiting absorption principles for the Navier equation in elasticity},
Ann. Sc. Norm. Super. Pisa Cl. Sci. (5), \textbf{11} (2012), no. 4, 817--842.
%
\bibitem{BFPRV2012}
\bysame,
 \emph{A Born approximation for live loads in Navier elasticity},
SIAM J. Math. Anal., \textbf{44} (2012), no. 4, 2824--2846.
%
\bibitem{BFPRV2018}
\bysame,
\emph{Uniqueness for inverse elastic medium problems},
SIAM J. Math. Anal., \textbf{50} (2018), no. 4, 3939--3962.
%
\bibitem{BD2013}
A.~Bchatnia, and M.~Daoulatli,
\emph{Behavior of the energy for Lamé systems in bounded domains with nonlinear damping and external force}, 
Electron J Differential Equations, \textbf{2013} (2013), 1--17.
%
\bibitem{BFV2014}
E.~Beretta, E.~Francini, and S.~Vessella, \emph{Uniqueness and Lipschitz stability for the identification of Lamé parameters from boundary measurements}, 
Inverse Problems Imaging, \textbf{8}
(2014), 611–644.
%
\bibitem{BG2014}
A.~Bchatnia, and A.~Guesmia, 
\emph{Well-posedness and asymptotic stability for the Lamé system with infinite memories in a bounded domain}, 
Mathematical Control and Related Fields, \textbf{4} (2014), 451--463.
%
\bibitem{BG2019}
A.~Benaissa, and S.~Gaouar, 
\emph{Asymptotic stability for the Lamé system with fractional boundary damping}, 
Comput. Math. Appl., \textbf{77} (2019), 1331--1346.
%
\bibitem{B-R2020}
L.E.~Bocanegra-Rodr\'iguez, M.A.~Jorge Silva, T.F.~Ma, and P.N.~Seminario-Huertas, 
\emph{Longtime dynamics of a semilinear Lamé system},
arXiv:2003.03646v1[math.AP] (2020).
%
\bibitem{BPST2003}
N.~Burq, F.~Planchon, J.G.~Stalker, and A.S. Tahvildar-Zadeh,
\emph{Strichartz estimates for the wave and Schrödinger equations with the inverse-square potential},
J. Funct. Anal., \textbf{203} (2003), no. 2, 519--549.
%
\bibitem{BPST2004}
\bysame,
\emph{Strichartz estimates for the wave and Schrödinger equations with potentials of critical decay},
Indiana Univ. Math. J., \textbf{53} (2004), no. 6, 1665--1680.
%
\bibitem{CLLZ2021}
P.~Caro, R.-Y.~Lai, Y.-H.~Lin, T.~Zhou,
\emph{Boundary determination of electromagnetic and Lamé parameters with corrupted data}, 
to appear in Inverse Problems \& Imaging (2021).
%	
\bibitem{CCF2021}
B.~Cassano, L.~Cossetti, and L.~Fanelli, \emph{Eigenvalue bounds and spectral stability of Lamé operators with complex potentials}, J. Diff. Eq. \textbf{298} (2021), 528--559.
%
\bibitem{Cossetti2017}
L.~Cossetti,
\emph{Uniform resolvent estimates and absence of eigenvalues for Lamé operators with subordinated complex potentials}
	J. Math. Anal. Appl., \textbf{1} (2017), 336--360.
%
\bibitem{Cossetti2019}
\bysame, 
\emph{Bounds on eigenvalues of perturbed Lamé operators with complex potentials}, 
arXiv:1904.08445[math.SP] (2019).
%
\bibitem{DN2002}
E.B.~Davies, and J.~Nath, 
\emph{Schrödinger operators with slowly decaying potentials}, 
J. Comput. Appl. Math., \textbf{148} (2002), 1--28.
%
\bibitem{ER2002}
G.~Eskin, and J.~Ralston, 
\emph{On the inverse boundary value problem for linear isotropic elasticity}, 
Inverse Problems, \textbf{18} (2002), 907--921.
%
\bibitem{FZ2016}
L.D.~Faddeev, and V.E.~Zakharov, 
\emph{Korteweg De Vries Equation: A Completely Integrable Hamiltonian System}, Fifty Years of Mathematical Physics (2016), 277--284.
%
\bibitem{FKV2018}
	L.~Fanelli, D.~Krej\v{c}i\v{r}\'ik, and L.~Vega, 
\emph{Spectral stability of Schr\"odinger operators with subordinated complex potentials}, 
	J. Spectr. Theory, \textbf{8} (2018), 575-604.
%	
\bibitem{F_K_V2}
	\bysame,
	\emph{Absence of eigenvalues of two-dimensional magnetic Schr\"odinger operators},
	J. Funct. Anal. \textbf{275} (2018), no. 9, 2453-2472.
%	
\bibitem{IY2015}
Y.~Imanuvilov, and M.~Yamamoto, 
\emph{Global uniqueness in inverse boundary value problems for the Navier Stokes equations and Lamé system in two dimensions},
Inverse Problems, \textbf{31} (2015), 121--142.
%
\bibitem{KKLS2021}
S.~Kim, Y.~Kwon, S.~Lee, and I.~Seo,
\emph{Strichartz and uniform Sobolev inequalities for the elastic wave equation},
arXiv:2102.04684v1[math.AP] (2021).
%
\bibitem{KKS2021}
S.~Kim, Y.~Kwon, and I.~Seo, 
\emph{Strichartz estimates and local regularity for the elastic wave equation with singular potentials}, 
Discrete \& Continuous Dynamical Systems \textbf{41} (2021), no. 4, 1897--1911.
%
\bibitem{KL2020}
Y.~Kwon, and S.~Lee, 
\emph{Sharp resolvent estimates outside of the uniform boundedness range}, Commun. Math. Phys., \textbf{374} (2020), no. 3, 1417--1467.
%
\bibitem{KLS2021}
Y.~Kwon, S.~Lee, and I.~Seo,
\emph{Resolvent estimates for the Lamé operator and failure of Carleman estimates}, to appear in J. Fourier Anal. Appl.
%
\bibitem{KK2020}
D.~Krej\v{c}i\v{c}\'ik, and T.~Kurimaiov\'a, 
\emph{From Lieb–Thirring Inequalities to Spectral Enclosures for the Damped Wave Equation}, 
Integr. Equ. Oper. Theory \textbf{92} (2020), 47.
%
\bibitem{Laptev2012}
A.~Laptev, 
\emph{Spectral inequalities for Partial Differential Equations and their applications}, 
AMS/IP Stud. Adv. Math \textbf{51} (2012), 629--643.
%
\bibitem{LW2000}
A.~Laptev, and T.~Weidl, 
\emph{Recent results on Lieb-Thirring inequalities}, 
Journ\'ees \'Equations aux d\'eriv\'ees partielles (2000), 1--14.
%
\bibitem{LN2017}
Y.-H.~Lin, and G.~Nakamura,
\emph{Boundary determination of the Lamé moduli for the isotropic elasticity system},
Inverse Problems, \textbf{33} (2017), 125004.
%
\bibitem{Mustafa1}
M.I.~Mustafa, 
\emph{Optimal decay rates for the viscoelastic wave equation}, Math Methods Appl Sci., \textbf{41} (2018), 192--204.
%
\bibitem{Mustafa2}
M.I.~Mustafa,
\emph{General decay result for nonlinear viscoelastic equations}, 
J Math Anal Appl., \textbf{457} (2018), 134--152.
%
\bibitem{NU1994}
G.~Nakamura, and G.~Uhlmann, 
\emph{Global uniqueness for an inverse boundary value problem
arising in elasticity}, 
Invent. Math., \textbf{118} (1994), 457--474.
%
\bibitem{RSIV}
M.~Reed, and B.~Simon, 
\emph{Methods in mathematical physics, vol. IV: Analysis of operators}, 
Academic Press, 1978.
%
\bibitem{Tartar1997}
L.~Tartar,
\emph{On the characterization of traces of a Sobolev space used for Maxwell's equation},
in: A.-Y. Le Roux (Ed.), proceedings of a Meeting held in Bordeaux, in Honour of Michel Artola, November, 1997.
%\bibitem{AFT2003}
%B.~Alziary, J.~Fleckinger-Pellé, and P.~Tak\'a\v c, 
%\emph{Eigenfunctions and Hardy inequalities for a magnetic Schrodinger operator in $\R^2$}, 
%Math. Meth. Appl. Sci. \textbf{26} (2003), 1093--1136.
%%
%\bibitem{BEL}
%A.~Balinsky, W.~Evans, and R. Lewis,  
%\emph{The analysis and geometry of Hardy’s inequality},
%Springer (2015). 
%%
%\bibitem{BLS2004}	
%A.~Balinsky, A.~Laptev, and A.~Sobolev, 
%\emph{Generalized Hardy inequality for the magnetic Dirichlet forms},
% J. Stat. Phys. \textbf{116} (2004), 507--521.
% %
% \bibitem{Beckner2008}
% W. Beckner, Weighted inequalities and Stein-Weiss potentials, Forum Math. \textbf{20} (2008), no. 4, 587–606.
% %
% \bibitem{Bennett1989}
% D. M.~Bennett,
% \emph{An extension of Rellich's inequality},
% Proc. Amer. Math. Soc. \textbf{106} (1989), 987--993.
% %
% \bibitem{BGR21}
% E.~Berchio, D.~Ganguly, and P.~Roychowdhury,
% \emph{Hardy-Rellich and second order Poincar\'e identities on the hyperbolic space via Bessel pairs}, arXiv:2106.03166[math.AP].
% %
% \bibitem{CFKP2021}
% B.~Cassano, V.~Franceschi, D.~Krej\v{c}i\v{r}\'ik, and D.~Prandi,
% \emph{Horizontal magnetic fields in the Heisenberg group}, (2021)
% %
% \bibitem{CP2018}
% B. Cassano, and F. Pizzichillo, 
% \emph{Self-adjoint extensions for the Dirac operator with Coulomb-type spherically symmetric potentials}, 
% Letters in Mathematical Physics, \textbf{108} (2018), no. 12, 2635--2667.
% %
%\bibitem{CPV2020}
%	B. Cassano, F. Pizzichillo and L. Vega,
%	\emph{A Hardy-type inequality and some spectral characterizations for the Dirac–Coulomb operator},
%	Rev. Mat. Complut. \textbf{33} (2020), 1-18.  
% %
% \bibitem{Cazacu2019}
% C.~Cazacu, 
% \emph{A new proof of the Hardy–Rellich inequality in any dimension}, 
% Proceedings of the Royal Society of Edinburgh: Section A Mathematics, \textbf{150} (2020), no. 6, 2894--2904.
% %
% \bibitem{CK2016}
% C. Cazacu and D. Krej\v ci\v r\'ik, 
% \emph{The Hardy inequality and the heat equation with magnetic field in any dimension}, 
% Comm. Partial Differential Equations \textbf{41} (2016), 1056--1088.	
%%
%\bibitem{CT2010}
%Y.~Colin de Verdière, and F. Truc,
%\emph{Confining quantum particles with a purely magnetic field}, Annales de l'Institut Fourier,  \textbf{60} (2010), no. 7, 2333--2356.
%%
%\bibitem{DH1998}
%E. B.~Davies, and A. M.~Hinz,
%\emph{Explicit constants for Relluch inequalities in $L^p(\Omega)$},
%Math. Z. \textbf{227} (1998), 511-523.
%%
%\bibitem{EP2014}
%T.~Ekholm, and F.~Portmann, 
%\emph{A magnetic contribution to the Hardy inequality}, 
%J. Math. Phys. \textbf{55} (2014), 022101.
%%
%\bibitem{EL2005}
%W.~D.~Evans, and R.~T.~Lewis,
%\emph{On the Rellich inequality with magnetic potentials},
%Math. Z., \textbf{251} (2005), no. 2, 267--284.
%%
%\bibitem{Fanelli2017}
%L.~Fanelli,
%\emph{Spherical Schrödinger Hamiltonians: Spectral analysis and time decay},
%In: Michelangeli A., Dell'Antonio G. (eds) Advances in Quantum Mechanics. Springer INdAM Series, vol 18. Springer, Cham. 2017.
%%
%\bibitem{FFFP2013}
%L.~Fanelli, V.~Felli, M. A.~Fontelos, and A. Primo, 
%\emph{Time decay of scaling critical electromagnetic Schrödinger flows}, 
%Commun. Math. Phys. \textbf{324} (2013), 1033–-1067.
%%
%\bibitem{FKLV}
%L.~Fanelli, D.~Krej\v ci\v r\'ik, A.~Laptev, and L.~Vega,
%\emph{On the improvement of the Hardy inequality due to singular magnetic fields}, 
%Communications in Partial Differential Equations, \textbf{45} (2020), no. 9, 1202--1212.
%%
%\bibitem{FFT2011}
%V.~Felli, A.~Ferrero, and S.~Terracini,
%\emph{Asymptotic behavior of solutions to Schrödinger equations near an isolated singularity of the electromagnetic potential}, Journal of the European Mathematical Society \textbf{013} (2011), no. 1, 119--174. 
%%
%\bibitem{FMT2007}
%V.~Felli, E.~Marchini, and S.~Terracini, 
%\emph{On the behavior of solutions to Schrödinger equations with dipole-type potentials near the singularity}, 
%Discrete and Continuous Dynamical Systems \textbf{21} (2007), 91-119.
%%
%\bibitem{FH2009}
%S.~Fournais, and B.~Helffer,  
%\emph{Diamagnetism}, 
%Spectral Methods in Surface Superconductivity. Progress in Nonlinear Differential Equations and Their Applications, vol. 77, Boston: Birkhäuser (2009), 19--30.
%%
%\bibitem{GM2011}
%N.~Ghoussoub, and A.~Moradifam, 
%\emph{Bessel pairs and optimal Hardy and Hardy-Rellich inequalities}, 
%Math. Ann. \textbf{349} (2011), no. 1, 1--57.
%%
%\bibitem{Hardy20}
%G.~H.~Hardy, 
%\emph{Note on a theorem of Hilbert}, 
%Math. Zeit. \textbf{6} (1920), 314--317
%%
%\bibitem{IKL}
%O.~Ibrogimov, D.~Krej\v ci\v r\'ik, and A. Laptev,
%\emph{Sharp bounds for eigenvalues of biharmonic operators with complex potentials in low dimensions},
%Mathematische Nachrichten (2021), 1--17.
%%
%\bibitem{Kato1972}
%T.~Kato, 
%\emph{Schrödinger operators with singular potentials}, 
%Isr. J. Math. \textbf{13} (1972), (1–2), 135--148.
%%
%\bibitem{Kovarik2011}
%H.~Kova\v r\'ik, 
%\emph{Eigenvalue bounds for two-dimensional magnetic Schrödinger operators}, 
%J. Spectr. Theory \textbf{1} (2011), 363--387.
%%
%\bibitem{Krejcirik2007}
%D.~Krej\v ci\v r\'ik,
%\emph{Twisting versus bending in quantum waveguides},
%Analysis on Graphs and its Applications,
%Cambridge, 2007 (P.Exner et al., ed.), Proc. Sympos. Pure Math., vol. 77, Amer. Math. Soc., Providence, RI, 2008, pp. 617-636.
%%
%\bibitem{Krejcirik2013}
%D.~Krej\v ci\v r\'ik,
%\emph{The improved decay rate for the heat semigroup with local magnetic field in
%the plane}, 
%Calc. Var. Partial Differ. Equ. \textbf{47} (2013), 207--226.
%%
%\bibitem{Kuwabara1982}
%R.~Kuwabara, 
%\emph{On spectra of the Laplacian on vector bundles} 
%Jour. Math. Tokushima Univ. \textbf{16} (1982), 1--23.
%%
%\bibitem{Kuwabara1988}
%R.~Kuwabara. 
%\emph{Spectrum of the Schrödinger operator on a line bundle over complex projective spaces}, 
%Tohoku Math. Jour. \textbf{40} (1988), 199--211.
%%
%\bibitem{LW1999}
%A.~Laptev, and T.~Weidl, 
%\emph{Hardy inequalities for magnetic Dirichlet forms}, 
%Oper. Theory Adv. Appl. \textbf{108} (1999), 299--305.
%%
%\bibitem{Lewis1982}
%R.T.~Lewis,
%\emph{Singular elliptic operators of second order with purely discrete spectra},
%Trans. Amer. Math. Soc. \textbf{271} (1982), 653-666.
%%
%\bibitem{Li_Lo}
%	E. H. Lieb and M. Loss,
%	\emph{Analysis}, 
%	Second Edition, American Mathematical Society, Providence, Rhode Island (2001).
%%	
%\bibitem{Rellich53}
%F.~Rellich,
%\emph{Halbbeschränkte Differentialoperatoren höherer Ordnung}. 
%In Proceedings of the International Congress of Mathematicians, 1954, Amsterdam, vol. III, 243--250. Erven P. Noordhoof N.V., Groningen, 1956. 
%%
%\bibitem{RB1969}
%F.~Rellich, and J.~Berkowitz,
%\emph{Perturbation theory of eigenvalue problems},
%Gordon and Breach Science Publishers, New York, London, Paris, 1969.
%%
%%\bibitem{Schmincke1972}
%%U.-W.~Schmincke,
%%\emph{Essential selfadjointness of a Schrödinger operator with strongly singular potential},
%%Math. Z. \textbf{124} (1972), 47--50.
%%
%\bibitem{Solomyak1994}
%M.~Solomyak,
%\emph{A remark on the Hardy inequalities},
%Integral Equations Operator Theory \textbf{19} (1994), 120--124.
%%
%\bibitem{TZ2007}
%A.~Tertikas, and N.~B.~Zographopoulos, 
%\emph{Best constants in the Hardy-Rellich inequalities and related improvements}, 
%Adv. Math. \textbf{206} (2007), no. 2, 407--459.
%%
%\bibitem{Thomas2007}
%J.~C.~Thomas, 
%\emph{Some Problems Associated with Sum and Integral Inequalities}, 
%Ph.D. thesis, Cardiff University, Wales (2007).
%%
%\bibitem{Torki_Hamza1989}
%N.~Torki-Hamza, 
%\emph{Stabilité des valeurs propres et champ magnétique sur une variété riemannienne et sur un graphe}, 
%PhD thesis, Grenoble University (1989).
%%
%\bibitem{VZ2000}
%J. L.~V\'azquez, and E.~Zuazua, 
%\emph{The Hardy inequality and the asymptotic behaviour of the heat equation with an inverse-square potential},  J. Funct. Anal.  \textbf{173} (2000), 103--153.
%%
%\bibitem{Weidl1999}
%T.~Weidl, 
%\emph{A remark on Hardy type inequalities for critical Schrödiinger operators with magnetic fields}, 
%Oper. Theory Adv. Appl. \textbf{110} (1999), 
%345--352.
%%
%\bibitem{Yafaev1999}
%D.~Yafaev,
%\emph{Sharp constants in the Hardy-Rellich inequalities},
%J. Funct. Anal. \textbf{168} (1999), 121--144.
%
%\bibitem{A_A_D}
%	A. A. Abramov, A. Aslanyan and E. B. Davies, 
%	\emph{Bounds on complex eigenvalues and resonances}, J. Phys. A \textbf{34} (2001), 57-72.
%%	
%%\bibitem{Agmon}
%%	S. Agmon,
%%	\emph{Spectral properties of Schr\"odinger operators and scattering theory},
%%	Ann. Scuola Norm. Sup. Pisa Cl. Sci. (4) \textbf{2} (1975), no. 2, 151-218.
%%	
%\bibitem{B_W}
%	R. Ba\~nuelos and G. Wang, 
%	\emph{Sharp inequalities for martingales with applications to the Beurling-Ahlfors and Riesz transforms}, 
%	Duke Math. J. \textbf{80} (1995), no. 3, 575-600.
%%	
%\bibitem{B_B_R_V}
%	J. A. Barcel\'o, J. M. Bennet, A. Ruiz and M. C. Vilela,
%	\emph{Local smoothing for Kato potentials in three dimensions},
%	Math. Nachr. \textbf{282} (2009), 1391-1405.
%%	

%%
%\bibitem{B_T_G20}
%	J. Behrndt, A.F.M. ter Elst and F. Gesztesy,
%	\emph{The generalized Birman-Schwinger Principle},
%	arXiv: 2005.01195v4 [math.SP] (2020).
%%
%%\bibitem{Birman}
%	%M. Sh. Birman, 
%	%\emph{The spectrum of singular boundary problems},
%	%Math. Sb. 55 (1966), 125-174; English transl., Amer. Math. Soc. Trans (2) 53 (1966), 23-80.
%	%
%	
%%\bibitem{Birman2}
%%	M. Sh. Birman,
%%	\emph{Discrete spectrum in the gaps of a continuous one for perturbations with large coupling constant},
%%	in Estimates and Asymptotics for Discrete Spectra of Integral and Differential Equations (Leningrad, 1989-90),
%%	Adv. Soviet Math. \textbf{7}, 
%%	Amer. Math. Soc., Providence (1991), 57-73.
%%
%%\bibitem{Bogli}
%%	S. B\"ogli,
%%	\emph{Schr\"odinger operator with non-zero accumulation points of complex eigenvalues},
%%	Comm. Math. Phys. \textbf{352} (2017), no. 2, 629-639.
%%
%\bibitem{BK08}
%	D. Borisov and D. Krej\v ci\v r\' ik,
%	\emph{$\mathcal{PT}$-symmetric wave guides},
%	Integral Equations Operator Theory \textbf{62} (2008), no. 4, 489-515.
%%
%%\bibitem{B_O}
%%	V. Bruneau and E. M. Ouhabaz, 
%%	\emph{Lieb-Thirring estimates for non self-adjoint
%%Schr\"odinger operators},
%%	J. Math. Phys. \textbf{49} (2008), no. 9, 093504, 10 pp.
%%
%\bibitem{Ca_Zy}
%	A.P. Calder\'on and A. Zygmund,
%	\emph{On singular integrals}, 
%	Amer. J. Math. \textbf{78} (1956), 289-309.
%%	
%\bibitem{C_F_L}
%	E. A. Carlen, R. L. Frank and E. H. Lieb,
%	\emph{Stability estimates for the lowest eigenvalue of a Schr\"odinger operator},
%	Geom. Funct. Anal. \textbf{24} (2014), no. 1, 63-84.
%%	

%%
%\bibitem{Ca_Ib_Kr_St19}
%	B. Cassano, O. O. Ibrogimov, D. Krej\v{c}i\v{r}\'ik and F. \v{S}tampach, 
%	\emph{Location of eigenvalues of non-self-adjoint discrete Dirac operators}, 
%	Ann. Henri Poincaré \textbf{21} (2020), 2193-2217.
%%
%%\bibitem{C_H_L}
%%	S. Chanillo, B. Helffer and A. Laptev, 
%%	\emph{Nonlinear eigenvalues and analytic hypoellipticity},
%%	J. Funct. Anal. \textbf{209} (2004), no. 2, 425-433.
%%	
%%\bibitem{Ch_Sa}
%%	S. Chanillo and E. Sawyer,
%%	\emph{Unique continuation for $\Delta + v$ and the C. Fefferman-Phong class},
%%	Trans. Amer. Math. Soc. \textbf{318} (1990), no. 1, 275-300.
%
%\bibitem{Ch_Fr}
%	F. Chiarenza and M. Frasca,
%	\emph{A remark on a paper by C. Fefferman},
%	Proc. Amer. Math. Soc. \textbf{108} (1990), 407-409.
%%
%%\bibitem{Ch_Ru}
%%	F. Chiarenza and A. Ruiz,
%%	\emph{Uniform $L^2$- weighted Sobolev inequalities},
%%	Proc. Amer. Math. Soc. \textbf{112} (1991), no. 1, 53-64.
%%		
%\bibitem{C_F}
%	R. R. Coifman and C. Fefferman,
%		\emph{Weighted norm inequalities for maximal functions and singular integrals},
%		Studia Math \textbf{51} (1974), 241-250.
%%		
%\bibitem{Co17}
%	L. Cossetti,
%	\emph{Uniform resolvent estimates and absence of eigenvalues for Lamé operators with subordinated complex potentials}
%	J. Math. Anal. Appl. \textbf{1} (2017), 336-360.
%%
%\bibitem{Co19}
%	\bysame,
%	\emph{Bounds on eigenvalues of perturbed Lamé operators with complex potentials},
%	arXiv:1904.08445 [math. SP] (2019).
%%	
%\bibitem{CM20}
%	L. Cossetti and R. Mandel,
%	\emph{A limiting absorption principle for Helmholtz systems and time-harmonic isotropic Maxwell's equations},
%	arXiv:2009.05087 [math.AP] (2020).
%%
%\bibitem{Cu14}
%	J.-C. Cuenin, 
%	\emph{Estimates on complex eigenvalues for Dirac operators on the half-line}, 
%	Integral Equations and Operator Theory \textbf{79} (2014), no.3, 377-388.
%%
%\bibitem{Cuenin}
%		J.-C. Cuenin,
%		\emph{Eigenvalue bounds for Dirac and Fractional Schr\"odinger operators with complex potentials},
%		J. Funct. Anal. \textbf{272} (2017), 2987-3018.
%%
%\bibitem{Cu19}
%	\bysame,
%	\emph{Eigenvalue estimates for bilayer graphene},
%	Ann. Henri Poincaré \textbf{20} (2019), 1501-1516.
%%	
%\bibitem{Cu_La_Tr14}
%	J.-C. Cuenin, A. Laptev and C. Tretter, 
%	\emph{Eigenvalue estimates for non-selfadjoint Dirac operators on the real line}, 
%	Ann. Henri Poincaré \textbf{15} (2014), no. 4, 707–736.
%%		
%\bibitem{Cu_Si18}	
%	J.-C. Cuenin and P. Siegl, 
%	\emph{Eigenvalues of one-dimensional non-selfadjoint Dirac operators and applications}, 			
%	Letters in mathematical physics \textbf{108} (2018), no.7, 1757-1778.
%%
%%\bibitem{Cw}
%%	M. Cwikel, 
%%	\emph{Weak type estimates for singular values and the number of bound states of Schr\"odinger operators},
%%	Ann. Math. \textbf{106} (1977), 93-102.
%%
%\bibitem{Davies}
%	E.B. Davies, 
%	\emph{Non-self-adjoint differential operators},
%	Bull. London Math. Soc. \textbf{34} (2002), no. 5, 513-532.
%%
%\bibitem{D_H_K}
%	M. Demuth, M. Hansmann and G. Katriel,
%	\emph{On the discrete spectrum of non-selfadjoint operators},
%	J. Funct. Anal. \textbf{257} (2009), no. 9, 2742-2759.
%%	
%\bibitem{D_H_K13}
%	\bysame
%	\emph{Lieb-Thirring type inequalities for Schr\"odinger operators with a complex-valued potential},
%	Integral Equations Operator Theory \textbf{75} (2013), no. 1, 1-5.
%%
%\bibitem{DA_Fa_Sc20}
%	P. D'Ancona, L. Fanelli and N. M. Schiavone,
%	\emph{Eigenvalue bounds for non-selfadjoint Dirac operators}, 
%      arXiv:2006.02778 [math.SP] (2020).
%%
%\bibitem{Ed_Ev}
%	D. E. Edmunds and W. D. Evans, 
%	\emph{Spectral theory and differential operators.} 
%	Oxford Mathematical Monographs. Oxford Science Publications. The Clarendon Press, Oxford University Press, New York, 1987. 
%%
%\bibitem{Enblom16}
%	A. Enblom,
%  	\emph{Estimates for Eigenvalues of Schrödinger Operators with Complex-Valued Potentials},
%  Letters in Mathematical Physics \textbf{106} (2016), 197-220.
%%
%\bibitem{En18}
%	A. Enblom, 
%	\emph{Resolvent estimates and bounds on eigenvalues for Dirac operators on the half-line},
%	 Journal of Physics A: Mathematical and Theoretical \textbf{51} (2018), no. 16, 165-203.
%%
%\bibitem{FK19}
%	L. Fanelli and D.Krej\v{c}i\v{r}\'ik,
%	\emph{Location of eigenvalues of three-dimensional non-self-adjoint Dirac operators},
%	Lett. Math. Phys. \textbf{109} (2019), no. 7, 1473–1485.
%%
%\bibitem{FKV18}
%	L. Fanelli, D. Krej\v{c}i\v{r}\'ik and L. Vega, 
%	\emph{Spectral stability of Schr\"odinger operators with subordinated complex potentials}, 
%	J. Spectr. Theory \textbf{8} (2018), 575-604.
%%	
%\bibitem{F_K_V2}
%	\bysame,
%	\emph{Absence of eigenvalues of two-dimensional magnetic Schr\"odinger operators},
%	J. Funct. Anal. \textbf{275} (2018), no. 9, 2453-2472.
%%	
%\bibitem{Fefferman1983}
%	C. Fefferman,
%	\emph{The uncertainty principle},
%	Bull. Amer. Math. Soc. \textbf{9} (1983), 129-206.
%%
%\bibitem{Fe_La_Sa19}
%	F. Ferrulli, A. Laptev and O. Safronov,
%	\emph{Eigenvalues of the bilayer graphene operator with a complex valued potential},
%	Analysis and Mathematical Physics \textbf{9} (2019), 1535-1546.
%%
%\bibitem{Frank}
%	R. L. Frank, 
%	\emph{Eigenvalue bounds for Schr\"odinger operators with complex potentials}, 
%	Bull. Lond. Math. Soc. \textbf{43} (2011), 745-750.
%%
%\bibitem{FrankIII}
%	\bysame,
%	\emph{Eigenvalue bounds for Schr\"odinger operators with complex potentials. III.},
%	Trans. Amer. Math. Soc. \textbf{370} (2018), no. 1, 219-240.
%%
%\bibitem{F_L_L_S}
%	R.L. Frank, A. Laptev, E. H. Lieb and R. Seiringer,
%	\emph{Lieb-Thirring inequalities for Schr\"odinger operators with complex-valued potentials},
%	Lett. Math. Phys. \textbf{77} (2006), no.3, 309-316.
%%
%\bibitem{F_L_S16}
%	R. L. Frank, A. Laptev and O. Safronov,
%	\emph{On the number of eigenavlues of Schr\"odinger operators with complex potentials}
%	J. London Math. Soc. \textbf{94} (2016), no. 2, 377-390.
%%
%\bibitem{Fr_Sa17}
%	R. L. Frank and J. Sabin, 
%	\emph{Restriction theorems for orthonormal functions, Strichartz inequalities, and uniform Sobolev estimates},  
%	Amer. J. Math. \textbf{139} (2017), no. 6, 1649-1691.
%%	
%\bibitem{FS17}
%	R. L. Frank and B. Simon,
%	\emph{Eigenvalue bounds for Schr\"odinger operators with complex potentials. II},
%	J. Spectr. Theory \textbf{7} (2017), 633-658. 
%%
%\bibitem{Ha_Kr20}
%	M. Hansmann and D. Krej\v{c}i\v{r}\'ik,
%    \emph{The abstract Birman-Schwinger principle and spectral stability}, 
%    arXiv: 2010.15102 [math.SP] (2020).
%%
%\bibitem{IKL19}
%	O. O. Ibrogimov, D. Krej\v{c}i\v{r}\'ik and A. Laptev,
%	\emph{Sharp bounds for eigenvalues of biharmonic operators with complex potentials in low dimensions}, 
%	arXiv:1903.01810 [math.SP] (2019).
%%
%\bibitem{Ib_St19}
%	O. O. Ibrogimov and F. \v{S}tampach, 
%	\emph{Spectral enclosures for non-self-adjoint discrete Schrödinger operators}, 
%	Integral Equations Operator Theory \textbf{91} (2019), no. 6, 53.
%%
%\bibitem{IJ2003}
%	A. D. Ionescu and D. Jerison,
%	\emph{On the absence of positive eigenvalues of Schrodinger operators with rough potentials},
%	Geom. Funct. Anal. \textbf{13} (2003), no. 5, 1029-1081.
%%
%%\bibitem{I_M}
%	%T. Iwaniec and G. Martin,
%	%\emph{Riesz transforms and related singular integrals} 
%	%Journal f\"ur die reine und angewandte Mathematik 473 (1995) 25-58.
%	%
%	\bibitem{Kato}
%		T. Kato, 
%		\emph{Perturbation theory for linear operators}, 
%		Die Grundlehren der mathematischen Wissenschaften, 132. Springer-Verlag, New York, New York, 1966.
%	%
%\bibitem{Keller}
%	J. B. Keller, 
%	\emph{Lower bounds and isoperimetric inequalities for eigenvalues of the Schr\"odinger equation},
%	J. Math. Phys. \textbf{2} (1961), 262-266.
%%	
%\bibitem{Ke_Sa1986}
%	R. Kerman and E. Sawyer, 
%	\emph{The trace inequality and eigenvalue estimates for Schrödinger operators}, 
%	Ann. Inst. Fourier (Grenoble) \textbf{36} (1986), 207-228.
%%
%\bibitem{K_R_S}
%	C. E. Kenig, A. Ruiz and C. D. Sogge,
%	\emph{Uniform Sobolev inequalities and unique continuation for second order constant coefficient differential operators},
%	Duke Math. J. \textbf{55} (1987), no.2, 329-347.
%%
%\bibitem{KT2002}
%	H. Koch and D. Tataru,
%	\emph{Sharp counterexamples in unique continuation for second order elliptic equations},
%	 J. reine angew. Math. \textbf{542} (2002), 133-146.
%%
%\bibitem{KT2006}
%	\bysame, 
%	\emph{Carleman estimates and absence of embedded eigenvalues}, 
%	Comm. Math. Phys. \textbf{267} (2006), no. 2, 419-449
%%
%\bibitem{Kr_Ku20}
%	D. Krej\v{c}i\v{r}\'ik and T. Kurimaiov\'a,
%	\emph{From Lieb–Thirring inequalities to spectral enclosures for the damped wave equation},
%	arXiv:2008.05176 [math.SP] (2020).
%%
%\bibitem{Kupradze}
%	V. D. Kupradze, 
%	\emph{Three-Dimensional Problems of Elasticity and Thermoelasticity}, 
%	North-Holland, Amsterdam (1979).
%%
%\bibitem{KL20}
%	Y. Kwon and S. Lee,
%	\emph{Sharp resolvent estimates outside of the uniform boundedness range},
%	Comm. Math. Phys. \textbf{374} (2020), no. 3, 1417-1467.
%%
%\bibitem{KLS20}
%	Y. Kwon, S. Lee and I. Seo,
%	\emph{Resolvent estimates for the Lamé operator and failure of Carleman estimates},
%	arXiv:1912.12620v2 [math.CA] (2020).
%%
%\bibitem{La_Li}
%	L. D. Landau, L. P. Pitaevskii, A. M. Kosevich and E. M. Lifshitz, 
%	\emph{Theory of Elasticity}, 
%	Third Edition, Butterworth-Heinemann, London (2012).
%%
%\bibitem{La_Sa}
%	A. Laptev and O. Safronov,
%	\emph{Eigenvalue estimates for Schr\"odinger operators with complex potentials},
%	Comm. Math. Phys. \textbf{292} (2009), no. 1, 29-54.
%%
%\bibitem{Lee_Seo}
%	Y. Lee and I. Seo,
%	\emph{A note on eigenvalue bounds for Schr\"odinger operators},
%	J. Math. Anal. Appl. \textbf{470} (2019), no.1, 340-347.
%%	
%%%\bibitem{L_S}
%%	%D. Levin and M. Solomyak, 
%%	%\emph{The Rozenblum-Lieb-Cwikel inequality
%%%for Markov generators}, 
%%	%J. Anal. Math. 71 (1997), 173-193.
%%	%
%%\bibitem{L1}
%%	E. H. Lieb, 
%%	\emph{The number of bound states of one-body Schr\"odinger operators and the Weyl problem},
%%	Geometry of the Laplace operator
%%	(Proc. Sympos. Pure Math., Univ. Hawaii, Honolulu, Hawaii, 1979), 241-252, Proc. Sympos. Pure Math., XXXVI, Amer. Math. Soc., Providence, R.I., 1980.
%%
%\bibitem{Li1983}
%	E. H. Lieb,
%	\emph{Sharp constant in the Hardy-Littlewood-Sobolev and related inequalities},
%	Ann. of Math. \textbf{2} (1983), no. 2, 349-374.
%%
%\bibitem{Li_Lo}
%	E. H. Lieb and M. Loss,
%	\emph{Analysis}, 
%	Second Edition, American Mathematical Society, Providence, Rhode Island (2001).
%%	
%%\bibitem{Li_Se}
%%	E. H. Lieb and R. Seiringer,
%%	\emph{The Stability of Matter in Quantum Mechanics},
%%	Cambridge University Press, 2010.
%%	
%\bibitem{L_T}
%	E. H. Lieb and W. Thirring, 
%	\emph{Inequalities for the moments of the
%eigenvalues of the Schr\"odinger Hamiltonian and their relation to
%Sobolev inequalities}, 
%Studies in Mathematical Physics, 
%Princeton University Press, Princeton (1976), 269-303.
%%
%%%\bibitem{L_Y}
%%	%P. Li and  S-T. Yau, 
%%	%\emph{On the Schr\"odinger equation and the eigenvalue problem}, 
%%	%Comm. Math. Phys. 88 (1983), 309-318.
%%
%\bibitem{Perez95}
%	C. P\'erez, 
%	\emph{Sharp $L^p$-weighted Sobolev inequalities}, 
%	Ann. Inst. Fourier (Grenoble) \textbf{45} (1995), no. 3, 809–824.
%%
%\bibitem{Petermichl}
%	S. Petermichl,
%	\emph{The sharp weighted bound for the Riesz transforms},
%	Proc. Amer. Math. Soc. \textbf{136} (2008), 1237-1249.
%%	
%%\bibitem{Ro1}
%%	G. V. Rozenblyum, 
%%	\emph{Distribution of the discrete spectrum of
%%singular differential operators},
%%	Sov. Math. Dokl. \textbf{13} (1972), 245-249.
%%	
%%\bibitem{Ro2}
%%	\bysame,
%%	\emph{Distribution of the discrete spectrum of singular differential operators}, 
%%	Soviet Math. (Iz. VUZ) \textbf{20} (1976), no. 1, 63-71.
%%
%%\bibitem{RV94}
%%	A. Ruiz and L.Vega,
%%	\emph{Local regularity of solutions to wave equations with time-dependent potentials},
%%	Duke Math. J. \textbf{76} (1994), no. 3, 913-940.
%%		
%\bibitem{Sa}
%	O. Safronov,
%	\emph{Estimates for eigenvalues of the Schr\"odinger operator with complex potential},
%	Bull. Lond. Math. Soc. 42 (2010), no. 3, 452-456.
%%	
%\bibitem{Sambou16}
%	D. Sambou, 
%	\emph{A criterion for the existence of non real eigenvalues for a Dirac operator}, 
%	New York J. Math \textbf{22} (2016), 469-500.
%%
%\bibitem{Seo}
%	I. Seo, 
%	\emph{From resolvent estimates to unique continuation for the Schrödinger equation}, 
%	Trans. Amer. Math. Soc. \textbf{368} (2016), 8755-8784.
%%
%\bibitem{Simon71}
%	B. Simon,		
%	\emph{Quantum mechanics for Hamiltonians defined as quadratic forms}.
%	Princeton Series in Physics. Princeton University Press, Princeton, N.J., 1971.
%%
%%\bibitem{Stepin}
%%	S.A. Stepin,  
%%	\emph{An estimate for the number of eigenvalues of the Schr\"odinger operator with complex potential}
%%	Sb. Math. \textbf{208} (2017), no. 1-2, 269-284.
%%	
%%\bibitem{Sylvester}
%%	J. Sylvester,
%%	\emph{An estimate for the free Helmholtz equation that scales},
%%	Inverse Probl. Imaging \textbf{3} (2009), no. 2, 333-351.
%%
%%\bibitem{Weidl}
%%	T. Weidl, 
%%	\emph{On the Lieb-Thirring constants
%%$L_{\gamma, 1}$ for $\gamma \geq \frac{1}{2}$}, 		
%%	Comm. Math. Phys. \textbf{178} (1996), no.1, 135-146.
\end{thebibliography}

\documentclass[a4paper,10pt,notitlepage,reqno]{article}
\usepackage[english]{babel}
\usepackage{amssymb,amsthm,amsmath} 
\usepackage{mathtools}
\usepackage{mathrsfs}
\usepackage{enumerate}
\usepackage{authblk}
\usepackage{cite}
\usepackage[symbol]{footmisc}
\usepackage{color}
\usepackage{geometry}
\usepackage[colorlinks=true, linkcolor=red, citecolor=blue, urlcolor=blue,hyperfootnotes=false]{hyperref}
\theoremstyle{plain} 
\newtheorem{theorem}{Theorem}[section]
\newtheorem{lemma}{Lemma}[section]

\newtheorem{corollary}{Corollary}[section]

\theoremstyle{definition}

\theoremstyle{remark}
\newtheorem{remark}{Remark}[section]

\DeclareMathOperator{\divergenza}{div}
\renewcommand{\div}{\divergenza}

\newcommand{\R}{\mathbb{R}}

\newcommand{\C}{\mathbb{C}}

\newcommand{\N}{\mathbb{N}}
\newcommand{\normeq}[1]{{\left\vert\kern-0.25ex\left\vert\kern-0.25ex\left\vert #1 
    \right\vert\kern-0.25ex\right\vert\kern-0.25ex\right\vert}}

{\left\lbrace\begin{array}{r@{\hspace{1mm}}ll}}%
{\end{array}\right.}

\title{\textbf{Spectral enclosures for the damped elastic wave equation}}
\author[1]{Biagio Cassano}
\author[2]{Lucrezia Cossetti} 
\author[3]{Luca Fanelli}		
\affil[1]{Dipartimento di Matematica e Fisica, Università degli Studi
  della Campania  ``Luigi Vanvitelli", Viale Lincoln 5, 81100 Caserta, Italy; biagio.cassano@unicampania.it}
\affil[2]{Fakult\"{a}t f\"{u}r Mathematik, Institut f\"{u}r Analysis, Karlsruher Institut f\"{u}r Technologie (KIT), Englerstra{\ss}e 2, 76131 Karlsruhe, Germany; lucrezia.cossetti@kit.edu}
\affil[3]{Ikerbasque $\&$ Departamento de Matematicas, Universidad del Pa\'is Vasco/Euskal Herriko Unibertsitatea (UPV/EHU), Aptdo. 644, 48080, Bilbao, Spain; luca.fanelli@ehu.es}
\begin{document}

\date{\small 17 August 2021}

\maketitle



\begin{abstract}
	\noindent
	In this paper we investigate spectral properties of the damped elastic wave equation. Deducing a correspondence between the eigenvalue problem of this model and the one of Lamé operators with non self-adjoint perturbations, we provide quantitative bounds on the location of the point spectrum in terms of suitable norms of the damping coefficient.
\end{abstract}

\section{Introduction}
This paper is concerned with the \emph{damped elastic wave equation}
\begin{equation}\label{eq:damped}
	u_{tt} + a(x)u_t-\Delta^\ast u=0, 
	\qquad (x,t)\in \R^d\times (0,\infty), 
\end{equation}
Here $a\colon \R^d \to \C^{d\times d}$ denotes the damping coefficient assumed to be a (possibly) non hermitian matrix. We shall make the standard assumption of a bounded damping, \emph{i.e.} $a\in L^\infty(\R^d)^d.$ The symbol $-\Delta^\ast$ is used to denote the Lamé operator of elasticity which is a matrix-valued differential operator acting, w.r.t. the spacial variable $x\in \R^d$ on smooth vector fields as
\begin{equation}\label{eq:Lame}
	-\Delta^\ast u=-\mu\Delta u -(\lambda+\mu)\nabla \div u,
	\qquad 
	u\in C^\infty_0(\R^d)^d:=C^\infty_0(\R^d;\C^d).
\end{equation}   
The material-dependent Lamé parameters $\lambda, \mu\in \R$ are assumed to satisfy the ellipticity condition
\begin{equation}\label{eq:ellipticity}
	\mu>0, 
	\quad
	\lambda+ \mu\geq 0.
\end{equation}

\noindent
It is customarily to write the second-order evolution system~\eqref{eq:damped} as a doubled first-order system introducing the vector field $U=(u, u_t)^{T}.$ Then~\eqref{eq:damped} can be rewritten as $U_t=A^\ast_a U,$ where $A_a^\ast$ is the $2d\times 2d$ matrix-valued \emph{damped elastic wave operator} defined as
\begin{equation}\label{eq:damped-operator}
	A_a^\ast:=
	\begin{pmatrix}
		0 &1\\
		\Delta^\ast & -a
	\end{pmatrix},
	\qquad \mathcal{D}(A_a^\ast):=H^2(\R^d)^d\times \dot{H}^1(\R^d)^d. 
\end{equation}   
The damped elastic wave equation~\eqref{eq:damped} and the corresponding damped operator~\eqref{eq:damped-operator} have attracted considerable attention in the last decades. In the constant coefficient case, namely $a(x)=\alpha,$ $\alpha>0,$ Bocanegra-Rodr\'iguez et al.~\cite{B-R2020} considered the longtime dynamics of this semilinear model in the presence of nonlinear structural forcing terms and external forces: they proved well-posedness \emph{\`a la} Hadamard and established the existence of finite dimensional global attractors together with the upper semicontinuity thereof. Energy decay results in relation with stability properties of solutions to this elastic model have been also deeply investigated. In~\cite{BD2013} Bchatnia and Daoulatli obtained a general energy decay estimate in a three dimensional bounded domain in the presence of localized nonlinear damping and an external force. By adding viscoelastic dissipation of memory type Bchatnia and Guesmia~\cite{BG2014} established a more general energy decay. Different viscoelastic dissipations have been considered in~\cite{Mustafa1,Mustafa2}.
Strong stability of Lamé systems with fractional order boundary damping were studied by Benaissa and Gaouar in~\cite{BG2019}.

For the undamped elastic wave equation, more commonly known as \emph{Navier equation}, a more varied bibliography is available. In~\cite{B_F-G_P-E_R_V} Barcel\'o et al.~proved uniform resolvent estimates (Limiting Absorption Principle) for this model. With this stationary tool at hands they also proved a priori averaged estimates for the corresponding Cauchy problem. The resolvent estimates in~\cite{B_F-G_P-E_R_V} were generalized in~\cite{Cossetti2019} and then improved in~\cite{KLS2021}, where a sharp result (analogous to the one available for the Laplacian~\cite{KL2020}) was proved. Surprisingly, differently from the Laplacian, in~\cite{KLS2021} the authors also showed the failure of uniform Sobolev and Carleman inequalities for the Lamé operator. In~\cite{KKLS2021} it was proved that if spacial lower-order  perturbations are replaced by temporal ones, \emph{i.e.} if one considers the damped equation, then those estimates become available again. In~\cite{B_F_R_V_V} the authors generalized the results in~\cite{B_F-G_P-E_R_V} proving Agmon-H\"ormander type estimates of the Navier equation when this is perturbed by small 0-th order matrix-valued potential. From these results Strichartz estimates for the evolution equation followed (in the same manner as for classical wave equation, see~\cite{BPST2003,BPST2004}). These Strichartz estimates were then generalized in~\cite{KKS2021,KKLS2021}. In particular in~\cite{KKLS2021} the endpoint case is deduced. 

The Navier equation got also attention of the inverse problem's community. In particular, inverse scattering was studied in~\cite{BFPRV2012,BFPRV2018}, whereas inverse boundary problems were considered in~\cite{NU1994,ER2002,BFV2014,IY2015,BFPRV2018}. Boundary determination of Lamé parameters has been studied in~\cite{CLLZ2021,LN2017,Tartar1997}.

\medskip
\noindent
In this paper we are interested in spectral properties of the damped elastic wave equation~\eqref{eq:damped}, or equivalently of the elastic wave operator~\eqref{eq:damped-operator}. More precisely, we aim at deducing quantitative bounds on the location of the point spectrum of $A_a^\ast$ in terms of suitable norms of the damping coefficient. In order to do that we establish a correspondence (see Lemma~\ref{lemma:correspondence}) between the eigenvalue problem associated to~\eqref{eq:damped-operator} and the one corresponding to suitable Lamé operators with non self-adjoint perturbations, that is operators of the form
\begin{equation}\label{eq:perturbed_Lame}
	-\Delta^\ast + V,
\end{equation} 
where $V$ denotes the operator of multiplication by a (possibly) non hermitian matrix-valued function $V\colon \R^d \to \C^{d\times d}$. 

The study of the spectrum of~\eqref{eq:perturbed_Lame} has already a bibliography. It is well known that the free Lamé operator $-\Delta^\ast$ is self-adjoint on $H^2(\R^d)^d$ and $\sigma(-\Delta^\ast)=\sigma_{\textup{ac}}(-\Delta^\ast)=[0,\infty).$ It is a natural question~\cite{Cossetti2017,Cossetti2019,CCF2021,
KLS2021} to ask whether and how the spectrum changes under 0th-order perturbations, \emph{i.e.} considering the operator~\eqref{eq:perturbed_Lame}. In~\cite{Cossetti2017}, adapting to the elasticity setting the method of multipliers developed for non self-adjoint Schrödinger operators in~\cite{FKV2018} (see also~\cite{F_K_V2} for similar problems on the plane), the author showed that the point spectrum of the perturbed Lamé operator~\eqref{eq:perturbed_Lame} remains empty (as in the free case) under suitable variational small perturbations (inverse-square Hardy potential with small coupling constant is covered). Later, in~\cite{CCF2021} we showed that full spectral stability, \emph{i.e.} $\sigma(-\Delta^\ast +V)=\sigma(-\Delta^\ast)=[0,\infty),$ can be proved in three dimensions $d=3$ under perturbations which satisfy a smallness condition of Hardy-type (see~\cite[Thm.~1.4]{CCF2021}). Focusing on the point spectrum only, if no stability can be proved a priori, an interesting question is related to provide quantitative bounds on the location in the complex plane of this part of the spectrum which, in the perturbed setting, is possibly no longer empty. In this direction, some preliminary result valid for the \emph{discrete} spectrum can be found in~\cite{Cossetti2019} (see also~\cite{KLS2021}). Later, these results have been extended in~\cite{CCF2021} to cover \emph{embedded} eigenvalues as well. More precisely in~\cite{CCF2021} the following result was proved.
\begin{theorem}[Thm.~1.1,~\cite{CCF2021}]\label{thm:Lp-result}
	Let $d\geq 2,$ $0<\gamma\leq 1/2$ if $d=2$ and $0\leq \gamma\leq 1/2$ if $d\geq 3$ and $V\in L^{\gamma + \frac{d}{2}}(\R^d;\C^{d\times d}).$ Then there exists a universal constant $c_{\gamma, d, \lambda, \mu}>0$ independent on $V$ such that
	\begin{equation}\label{eq:thesis}
		\sigma_\textup{p}(-\Delta^\ast + V)
		\subset \left\{z\in \C\colon |z|^\gamma \leq c_{\gamma, d, \lambda, \mu} \| V \|_{L^{\gamma + \frac{d}{2}}(\R^d)}^{\gamma + \frac{d}{2}}\right \}. 
	\end{equation}
\end{theorem}
\noindent
In the self-adjoint case, namely for real-valued perturbations, the result above holds for a larger class of indices $\gamma.$ More precisely, the following result holds true.
\begin{theorem}[Thm.~3.1,~\cite{Cossetti2019}]\label{thm:eigenvalue_bound_self}
	Let $d\geq 2,$ $\gamma>0$ if $d=2$ and $\gamma\geq 0$ if $d\geq3$ and $V\in L^{\gamma + \frac{d}{2}}(\R^d;\R).$ Then there exists a universal constant $c_{\gamma, d, \lambda, \mu}>0$ independent of $V$ such that any negative eigenvalue $z$ (if any) of the self-adjoint perturbed Lamé operator $-\Delta^\ast + VI_{\R^d}$ satisfies
	\begin{equation}\label{SA_bound_Lame}
		|z|^\gamma\leq 
		c_{\gamma, d, \lambda, \mu} \| V_- \|_{L^{\gamma + \frac{d}{2}}(\R^d)}^{\gamma + \frac{d}{2}},
	\end{equation}
	where $V_-$ is the negative part of $V,$ \emph{i.e.} $V_-(x):=\max\{-V(x),0\}.$
\end{theorem}
\noindent
Making use of Theorem~\ref{thm:Lp-result} and Theorem~\ref{thm:eigenvalue_bound_self} and the correspondence between the eigenvalue problem associated to the damped elastic wave operator and the one of the perturbed Lamé operator~\eqref{eq:perturbed_Lame} (see Lemma~\ref{lemma:correspondence} below) we shall prove the following two results valid in the self-adjoint and the non self-adjoint setting.
\begin{theorem}\label{thm:main-sa}
	Let $d\geq 2$ and assume $\gamma$ satisfies the hypotheses of Theorem~\ref{thm:eigenvalue_bound_self} and $a\in L^\infty(\R^d;\R).$ Then there exists a universal constant $c_{\gamma,d,\lambda, \mu}>0$ independent of the damping $a$ such that for any positive (respectively negative) eigenvalue $z$ of the damped elastic wave operator $A_a^\ast$ and $a_-\in L^{\gamma+ \frac{d}{2}}(\R^d)$ (respectively $a_+\in L^{\gamma+ \frac{d}{2}}(\R^d)$) satisfies
	\begin{equation}\label{eq:bound-sa}
		(\pm z)^{\gamma- \frac{d}{2}}\leq c_{\gamma,d,\lambda, \mu} \| a_{\mp} \|_{L^{\gamma + \frac{d}{2}}(\R^d)}^{\gamma + \frac{d}{2}},
	\end{equation}
\end{theorem}
\noindent
Setting $\gamma=d/2$ in~\eqref{eq:bound-sa}, the previous theorem provides sufficient condition on the size of the damping coefficient to guarantee absence of positive (respectively negative) eigenvalues.
\begin{corollary}\label{cor:absence-sa}
	If $d\geq 2$ and 
	\begin{equation*}
	c_{\frac{d}{2},d,\lambda, \mu} \|a_\mp\|_{L^d(\R^d)}^{d}<1,
	\end{equation*}
	then $A_a^\ast$ has no positive (respectively negative) eigenvalues.
\end{corollary}

\noindent
In the non self-adjoint setting we shall prove the following result. 
\begin{theorem}\label{thm:main-nsa}
	Let $d\geq 2$ and assume $\gamma$ satisfies the hypotheses of Theorem~\ref{thm:Lp-result} and $a\in L^\infty(\R^d;\C^{d\times d})$ is a (possibly) non hermitian matrix. Then there exists a universal constant $c_{\gamma,d,\lambda,\mu}>0$ independent of the damping $a$ such that
\begin{equation}\label{eq:bound-nsa}
	\sigma_{\textup{p}}(A_a^\ast)\subset 
	\Big\{
	z\in \C\colon |z|^{\gamma-\frac{d}{2}}
	\leq c_{\gamma, d,\lambda, \mu} \|a\|_{L^{\gamma+ \frac{d}{2}}}^{\gamma+ \frac{d}{2}}
	\Big\}.
\end{equation}	
\end{theorem}
\begin{remark}
	Notice that in the non self-adjoint case, due to the more restrictive class of indices for which Theorem~\ref{thm:Lp-result} is valid compared to Theorem~\ref{thm:eigenvalue_bound_self}, no analogous of Corollary~\ref{cor:absence-sa} holds true ($\gamma=d/2$ is not admissible).
\end{remark}

\medskip
\noindent
The main motivation behind our project relies on the following simple observation: the ellipticity condition~\eqref{eq:ellipticity} allows taking $\lambda+ \mu=0$ in the definition of the Lamé operator~\eqref{eq:Lame}. This choice turns the Lamé operator~\eqref{eq:Lame} into a vector Laplacian and consequently the damped elastic wave equation~\eqref{eq:damped} into a system of classical damped wave equations. For the (scalar) damped wave equation, results in the spirit of Theorem~\ref{thm:main-sa} and Theorem~\ref{thm:main-nsa} have been recently proved in~\cite{KK2020}. Thus, Theorem~\ref{thm:main-sa} and Theorem~\ref{thm:main-nsa} can be seen as a generalization of the results in~\cite[Thm.~1, Thm.~5 and Thm.~6]{KK2020} in the sense that they recover\footnote{the constants involved slightly differ due to the presence of the coefficient $\mu$ of the vector Laplacian and due to the vectorial form of the wave equation once $\lambda+ \mu=0$ in~\eqref{eq:damped}.} them when $\lambda+ \mu=0.$    

\medskip
\noindent
Theorem~\ref{thm:main-sa} and Theorem~\ref{thm:main-nsa} are not stated for $d=1,$ as a matter of fact the one dimensional case is rather special and it is treated separately. In $d=1$ the Lamé operator $-\Delta^\ast$ turns into a scalar differential operator, more precisely it is simply a multiple of the Laplacian
\begin{equation*}
	-\Delta^\ast
	:=-\mu \frac{d^2}{dx^2} -(\lambda + \mu) \frac{d^2}{dx^2}
	=-(\lambda + 2\mu) \frac{d^2}{dx^2}.
\end{equation*}
As a straightforward consequence of the celebrated result of Abramov, Aslanian and Davies for 1D-Schrödinger operators (see~\cite[Thm.~4]{AAD2001}), in~\cite{Cossetti2019} the following result for the one dimensional non self-adjoint Lamé operator was proved.
\begin{theorem}[Thm.~1.1,~\cite{Cossetti2019}]
\label{thm:1d}
	Let $d=1$ and $V\in L^1(\R;\C).$ Then
	\begin{equation*}
		\sigma_\textup{p}(-\Delta^\ast + V)\subset
		\Big\{z\in \C\colon |z|^{1/2}\leq \tfrac{1}{2\sqrt{\lambda+2\mu}} \|V\|_{L^1(\R)} \Big\}.
	\end{equation*} 
\end{theorem}
\begin{remark}
	We stress that Theorem 1.1 in~\cite{Cossetti2019} was stated only for eigenvalues outside the essential spectrum, namely for $z\in \C\setminus [0,\infty).$ Nevertheless, it is easy to show that embedded eigenvalues can be covered as well (see~\cite[Cor.~2.16]{DN2002}).
\end{remark}

\noindent
In the self-adjoint case, as an immediate consequence of the Lieb-Thirring inequalities (\hspace{-0.007cm}\cite{Laptev2012,LW2000}) valid for the Schrödinger operators, one has the following result.
\begin{theorem}\label{thm:1dLT}
	Let $d=1$ and $V_-\in L^1(\R;\R).$ Then
	\begin{equation}\label{eq:1dLT}
		\sigma_\textup{p}(-\Delta^\ast + V) \subset 
		\Big\{z\in \C\colon |z|^{1/2}\leq \tfrac{1}{2\sqrt{\lambda+2\mu}} \|V_-\|_{L^1(\R)} \Big\}.
	\end{equation}
\end{theorem}

\noindent
Theorem~\ref{thm:1d} and Theorem~\ref{thm:1dLT} together with Lemma~\ref{lemma:correspondence} below allow to deduce properties on the point spectrum of the one dimensional damped elastic wave operator $A_a^\ast.$
Differently from the higher dimensional setting, in $d=1$ Theorem~\ref{thm:1d} does not entail any quantitative bound on the location in the complex plane of the eigenvalues of $A_a^\ast,$ on the other hand it provides an explicit smallness condition on the size of the $L^1$-norm of the damping such that $A_a^\ast$ does not have eigenvalues. More precisely we have the following result.
\begin{theorem}\label{thm:1d-nsa}
	Let $d=1$ and $a\in L^1(\R;\C).$ If $\|a\|_{L^1(\R)}<2\sqrt{\lambda + 2\mu},$ then $\sigma_\textup{p}(A_a^\ast)=\varnothing.$ Moreover, the constant $2\sqrt{\lambda+ 2\mu}$ is optimal.
\end{theorem}

\noindent
In the self-adjoint situation it holds true a slightly different result compared to the ones introduced so far.
\begin{theorem}\label{thm:1d-sa}
Let $d=1$ and assume that $a$ is real-valued and satisfies 
\begin{equation}\label{eq:assumptions}
	\int_{\R} |x||a(x)|\, dx<\infty
	\qquad \text{and} \qquad
	\lim_{R\to \infty} \|a\|_{L^\infty(\R\setminus B_R(0))}=0.
\end{equation}
Let $z$ be a real eigenvalue of $A_a^\ast.$ If $z>0$ and $\int_{\R} a<-4\sqrt{\lambda + 2\mu}$ (or $z<0$ and $\int_{\R} a>4\sqrt{\lambda + 2\mu}$), then
\begin{equation*}
	|z|\geq (\lambda + 2\mu) \Bigg(\int_{\R} |x||a(x)|\, dx \Bigg)^{-1}.
\end{equation*}	
\end{theorem}
\noindent
Moreover the following quantitative bound on the location of eigenvalues holds.
\begin{theorem}\label{thm:1d-sa-bound}
	Let $d=1$ and assume that $a$ is real-valued and satisfies~\eqref{eq:assumptions}. Moreover, assume 
	\begin{equation*}
		|z|<(\lambda+ 2\mu)\Bigg(\int_{\R} |x||a(x)|\, dx \Bigg)^{-1}.
	\end{equation*}  
	If $z>0$ and $\int_\R a<0$ (respectively, $z<0$ and $\int_\R a>0$), then there exists exactly one $\alpha>0$ satisfying
	\begin{equation*}
 		2\Bigg(\int_\R a_-(x)\, dx\Bigg)^{-1}\leq \alpha \leq -4 \Bigg(\int_\R a(x)\, dx\Bigg)^{-1}
 		\quad \Bigg(\text{respectively, } 2\Bigg(\int_\R a_+(x)\, dx\Bigg)^{-1}\leq \alpha \leq 4 \Bigg(\int_\R a(x)\, dx\Bigg)^{-1} \Bigg) 
	\end{equation*}
	such that $z/\alpha$ is an eigenvalue of $A_{a}^\ast.$
\end{theorem}

\noindent
The rest of the paper is divided as follows. In the next Section we provide the proof of the preliminary Lemma~\ref{lemma:correspondence} establishing the correspondence between the eigenvalue problem associated to the damped elastic wave operator and the perturbed Lamé operator. Afterwards, in Section~\ref{sec:higher} we show the validity of Theorem~\ref{thm:main-sa} and Theorem~\ref{thm:main-nsa} which hold in higher dimension $d\geq 2.$ The one dimensional case, that is Theorem~\ref{thm:1d-nsa}-Theorem~\ref{thm:1d-sa-bound}, is treated separately in Section~\ref{sec:1d}.
\section{Proofs}
As a starting point we show how the eigenvalue problem associated to the damped elastic wave operator $A_a^\ast$ defined in~\eqref{eq:damped-operator} is related to the one of a perturbed Lamé operator of the form~\eqref{eq:perturbed_Lame}.
\begin{lemma}\label{lemma:correspondence}
Let $d\geq 1$ and assume $a\in L^\infty(\R^d;\C^{d\times d}).$ For every $z\in \C,$
\begin{equation*}
z\in \sigma_\textup{p}(A_a^\ast)
\quad  \Longleftrightarrow \quad 
-z^2\in \sigma_\textup{p}(-\Delta^\ast + za).
\end{equation*} 
\end{lemma}
\begin{proof}
	Assume $z\in \sigma_\textup{p}(A_a^\ast),$ then there exists a non-trivial $\Psi=(\psi_1,\psi_2)^T\in\mathcal{D}(A_a^\ast)$ such that $A_a^\ast \Psi=z\Psi.$ In other words, $\psi_1 \in H^2(\R^d)^d,$ $\psi_2\in \dot{H}^1(\R^d)^d$ and $\psi_2=z\psi_1,$ $\Delta^\ast \psi_1 -a \psi_2=z\psi_2.$ Plugging the first equation in the second one gives $-\Delta^\ast \psi_1 +za\psi_1=-z^2\psi_1.$ Since $\psi_1\neq 0,$ then $-z^2\in \sigma_\textup{p}(-\Delta^\ast +za).$ 
	Conversely, assume $-z^2\in  \sigma_\textup{p}(-\Delta^\ast +za),$ then there exists a non-trivial $\psi\in H^2(\R^d)^d$ such that $(-\Delta^\ast +za)\psi=-z^2 \psi.$ Defining $\Psi:=(\psi, z\psi)^T,$ then $\Psi \in \mathcal{D}(A_a^\ast)$ and 
		$(A_a^\ast \Psi)^T=(z \psi, \Delta^\ast \psi - za \psi)=z(\psi, z\psi)=z\Psi^T.$
	Therefore, $z\in \sigma_\textup{p}(A_a^\ast).$
 \end{proof}
 
 \begin{remark} \label{rmk:ev0}
From the validity of Lemma~\ref{lemma:correspondence}, one has that $0\notin \sigma_\textup{p}(A_a^\ast)$ as the spectrum of the unperturbed Lamé operator $-\Delta^\ast + 0 a=-\Delta^\ast$ is purely continuous.
 \end{remark}

\subsection{Higher dimensions $d\geq2:$ Proof of Theorem~\ref{thm:main-sa} and Theorem~\ref{thm:main-nsa}}\label{sec:higher}
With Lemma~\ref{lemma:correspondence} at hands we now show that Theorem~\ref{thm:main-sa} and Theorem~\ref{thm:main-nsa} are consequence of Theorem~\ref{thm:eigenvalue_bound_self} and Theorem~\ref{thm:Lp-result}, respectively. 
 
\begin{proof}[Proof of Theorem~\ref{thm:main-sa}]
 From Lemma~\ref{lemma:correspondence} we know that $z\in \sigma_\textup{p}(A_a^\ast)$ if and only if $-z^2\in \sigma(-\Delta^\ast + za).$ From Theorem~\ref{thm:eigenvalue_bound_self} there exists $c_{\gamma,d,\lambda, \mu}>0$ such that 
 \begin{equation}\label{eq:prel}
 	|z|^{2\gamma}\leq c_{\gamma,d,\lambda,\mu} \|(za)_-\|_{L^{\gamma + \frac{d}{2}}(\R^d)}^{L^{\gamma + \frac{d}{2}}},
 \end{equation}
 where $(za)_-$ is the negative part of $za,$ \emph{i.e.} $(za)_-= z a_+$ if $z\in (-\infty, 0)$ and $(za)_-=za_-$ if $z\in(0,\infty).$ Using this fact in~\eqref{eq:prel} and dividing both sides of~\eqref{eq:prel} by $|z|^{\gamma + d/2}$ ($z\neq 0,$ see Remark~\ref{rmk:ev0}) we obtain~\eqref{eq:bound-sa}.
\end{proof}
 
\medskip
\noindent 
Now we consider the non self-adjoint situation. 
\begin{proof}[Proof of Theorem~\ref{thm:main-nsa}]
The proof of Theorem~\ref{thm:main-nsa} is analogous to the one of Theorem~\ref{thm:main-sa}. Let $z\in \sigma_\textup{p}(A_a^\ast),$ then by Lemma~\ref{lemma:correspondence} $-z^2\in \sigma_\textup{p}(-\Delta^\ast + za).$ Using the eigenvalue bound~\eqref{eq:thesis} then one has
\begin{equation*}
|z|^{2\gamma}\leq c_{\gamma, d,\lambda, \mu} |z|^{\gamma + \frac{d}{2}} \|a\|_{L^{\gamma + \frac{d}{2}}(\R^d)}^{\gamma + \frac{d}{2}},
\end{equation*}
which gives~\eqref{eq:bound-nsa} and concludes the proof. 
\end{proof}

\subsection{1D: Proof of Theorem~\ref{thm:1d-nsa}, Theorem~\ref{thm:1d-sa} and Theorem~\ref{thm:1d-sa-bound}} \label{sec:1d}
We start considering the self-adjoint situation. Let $z\in \R$ and let $\{\lambda_n^\ast(za)\}_{n=1}^N$ denote the sequence of eigenvalues of $-\Delta^\ast + za,$ then the following preliminary lemma on the sum of the square root of the eigenvalues holds. 
\begin{lemma}
	Let $d=1.$ Then 
	\begin{equation}\label{eq:sum}
		\sum_{n=1}^N|\lambda_n^\ast(za)|^{1/2}\geq -\frac{z}{4\sqrt{\lambda+ 2\mu}} \int_\R a(x)\, dx.
	\end{equation}
	Moreover if   
		$\int_{\R} |x||a(x)|\, dx<\infty,$ then
	the following bound on the number $N$ of eigenvalues $\lambda_n^\ast(za)$
	\begin{equation}\label{eq:NLame}
		N\leq 1 + \frac{|z|}{\lambda + 2\mu} \int_{\R} |x||a(x)|\, dx
	\end{equation}
	holds.
\end{lemma}
\begin{proof}
	If $\lambda_n^\ast(za)$ is an eigenvalue of $-\Delta^\ast + za,$ then there exists $\psi\in H^2(\R)$ such that $-(\lambda + 2\mu)
	\Delta\psi+ za\psi=\lambda_n^\ast(za)\psi$ or, equivalently, 
	\begin{equation}\label{eq:Schro}
		\Big(-\Delta+ \frac{za}{\lambda + 2\mu}\Big)\psi=\frac{\lambda_n^\ast(za)}{\lambda + 2\mu}\psi.
	\end{equation}  
	Denoting by $\lambda_n(V)$ the eigenvalues of the Schrödinger operator $-\Delta+ V,$ then we conclude that $\lambda_n^\ast(za)$ is an eigenvalue of $-\Delta^\ast + za$ if and only if there exists $n\in \N$ such that $\lambda_n^\ast(za)$ is a multiple of an eigenvalue $\lambda_n(za/(\lambda + 2\mu))$ of the Schrödinger operator $-\Delta + za/(\lambda+ 2\mu),$ more precisely $\lambda_n(za/(\lambda + 2\mu))= \lambda_n^\ast(za)/(\lambda + 2\mu).$ In particular the number of eigenvalues coincides. The Buslaev-Faddeev-Zakharov trace formula (\emph{cf.}~\cite{FZ2016}) for 1D-Schrödinger operator $-\Delta + V$ states that
	\begin{equation*}
		\sum_{n=1}^N |\lambda_n(V)|^{1/2}\geq -\frac{1}{4} \int_{\R} V(x)\, dx,
	\end{equation*}
	this and the correspondence above give immediately~\eqref{eq:sum}.
	
	The Bargmann bound~\cite[Pb.~22]{RSIV} provides a control from above of the number of eigenvalues of the 1D-Schrödinger operator $-\Delta + V$ under the condition $\int_\R |x||V(x)\, dx<\infty.$ More precisely,
	\begin{equation}\label{eq:NSchro}
		N\leq 1 + \int_\R |x||V(x)|\, dx.
	\end{equation}
	Similarly as above (that is using the correspondence between eigenvalues of the Lamé operator $-\Delta^\ast + za$ and of the Schrödinger operator $-\Delta + za/(\lambda+ 2\mu)$) from~\eqref{eq:NSchro} one easily gets~\eqref{eq:NLame}. This concludes the proof.
\end{proof}
\begin{proof}[Proof of Theorem~\ref{thm:1d-sa}]
	Let $z$ be a real eigenvalue of $A_a^\ast,$ in order to prove Theorem~\ref{thm:1d-sa} we will show that if $|z|<(\lambda + 2\mu)\Big(\int_{\R} |x||a(x)|\, dx \Big)^{-1}$ then $\int_{\R} a\geq -4\sqrt{\lambda + 2\mu}$ for $z>0$ and $\int_{\R}a \leq 4\sqrt{\lambda+ 2\mu}$ for $z<0.$ First of all notice that~\eqref{eq:sum} is non-trivial only if $z\int_{\R} a(x)\, dx<0.$ This last condition, in particular is known to be a sufficient condition which guarantees that $\inf \sigma(-\Delta^\ast + za)<0.$ From the decay assumption~\eqref{eq:assumptions}, then it follows that $-\Delta^\ast + za$ posses at least one negative eigenvalue. From the upper bound~\eqref{eq:NLame} it follows that if $|z|<(\lambda + 2\mu)\Big(\int_{\R} |x||a(x)|\, dx \Big)^{-1}$ then $-\Delta^\ast + za$ has exactly one negative eigenvalue $\lambda_1^\ast(za).$ Thus, from~\eqref{eq:sum} and the correspondence in Lemma~\ref{lemma:correspondence} one has   
	\begin{equation}\label{eq:estimate1d}
		|z|=|\lambda_1(za)|^{1/2}\geq -\frac{z}{4\sqrt{\lambda + 2\mu}}\int_\R a(x)\, dx.
	\end{equation}
	This implies $\int_\R a(x)\, dx\geq -4\sqrt{\lambda+ 2\mu}$ for $z>0$ and $\int_\R a(x)\, dx\leq 4\sqrt{\lambda+ 2\mu}$ for $z<0.$
\end{proof}

\begin{proof}[Proof of Theorem~\ref{thm:1d-sa-bound}]
	From the hypotheses, as above, one has that $-\Delta^\ast + za$ posses exactly one negative eigenvalue. The Lieb-Thirring type bound~\eqref{eq:1dLT} in Theorem~\ref{thm:1dLT} and the estimate in~\eqref{eq:estimate1d} give
	\begin{multline*}
		-\frac{z}{4\sqrt{\lambda + 2\mu}} \int_{\R} a(x)\, dx\leq |\lambda_1(za)|^{1/2}\leq \frac{z}{2\sqrt{\lambda + 2\mu}} \int_{\R}a_-(x)\, dx,\\
		\qquad \Big(\text{respectively }
		-\frac{z}{4\sqrt{\lambda + 2\mu}} \int_{\R} a(x)\, dx\leq |\lambda_1(za)|^{1/2}\leq \frac{|z|}{2\sqrt{\lambda + 2\mu}} \int_{\R}a_+(x)\, dx \Big).
	\end{multline*}
	Using the correspondence in Lemma~\ref{lemma:correspondence} the result follows.
\end{proof}

\begin{proof}[Proof of Theorem~\ref{thm:1d-nsa}]
	If $z\in \C$ is an eigenvalue of $A_a^\ast,$ then from Lemma~\ref{lemma:correspondence} $-z^2\in \sigma_\textup{p}(-\Delta^\ast + za).$ Thus, from Theorem~\ref{thm:1d} we have
	\begin{equation*}
		|z|\leq \frac{1}{2\sqrt{\lambda+ 2\mu}} |z|\|a\|_{L^1(\R)}.
	\end{equation*}
	Dividing by $|z|,$ which cannot be zero (see Remark~\ref{rmk:ev0}), one has
		$1\leq \tfrac{1}{2\sqrt{\lambda+ 2\mu}} \|a\|_{L^1(\R)}.$
	If the $L^1$-norm of $a$ is small, namely if $\|a\|_{L^1(\R)}<2\sqrt{\lambda+ 2\mu},$ then we get a contradiction. Thus, $\sigma_\textup{p}(A_a^\ast)=\varnothing.$ The optimality of the result can be proved as in~\cite[Thm.~4]{KK2020}.
\end{proof}

\providecommand{\bysame}{\leavevmode\hbox to3em{\hrulefill}\thinspace}


\end{document}